If the ratio of final and initial intensities for 4% - solution of urea is 1/2 (length of beaker is 20 mm). Determine the ratio of intensities for 8%- solution of urea in the beaker of length l2 =30 mm.
1
Expert's answer
2020-10-23T11:54:19-0400

The intensity decreases with the square of the beaker length:


"\frac{I_f}{I_i}=\omega\frac{k}{d^2},\\\space\\
k=\frac{I_fd^2}{I_i\omega}."

For another beaker and solution:

"\frac{I_f'}{I_i'}=\frac{\omega'}{\omega }\frac{I_fd^2}{I_id'^2}=\frac{4}{9}"